OpenStudy (jkbo):

Please Help Use the following scenario to answer the question: A bag contains 5 green marbles, 4 red marbles, 7 blue marbles, and 9 yellow marbles. If 3 marbles are drawn from the bag one at a time and not replaced, what is the probability of drawing a blue, yellow, and red marble in that particular order? A.6/11 B.5/18 C.63/3450 D.1/48

OpenStudy (agent0smith):

5 green marbles, 4 red marbles, 7 blue marbles, and 9 yellow marbles. Find the total first... 5+4+7+9= 25 marbles 3 marbles are drawn from the bag one at a time and not replaced, what is the probability of drawing a blue, yellow, and red first: probability of drawing a blue... 7/25 second: probability of drawing a yellow: 9/24 (you've taken out one marble, leaving 24) third: probability of drawing red: 4/23 (you've taken out 2 marbles) Now multiply the probabilities together.

OpenStudy (jkbo):

which ones? @agent0smith

OpenStudy (agent0smith):

er, all of them.

OpenStudy (jkbo):

thank you

OpenStudy (agent0smith):

Hopefully you understand how i found those probabilities...

OpenStudy (jkbo):

I got 21/ 1180? @agent0smith

OpenStudy (jkbo):

wait sorry 21/1150

OpenStudy (agent0smith):

you can probably simplify it...\[\Large \frac{ 7 \times 9 \times 4 }{ 25 \times 24 \times 23 }\] if you write it as \[\Large \frac{ 7 \times 4 \times 9 }{ 25 \times 24 \times 23 }\] then you can already simplify the 4/24

OpenStudy (jkbo):

2/12

OpenStudy (jkbo):

@agent0smith

OpenStudy (agent0smith):

All i did was change the 9 into 3x3, and the 24 into 4x3x2 \[\Large \frac{ 7 \times 4 \times 3 \times 3 }{ 25 \times 4 \times 3 \times 2\times 23 }\]

OpenStudy (jkbo):

@agent0smith I got 252/ 13800

OpenStudy (jkbo):

do I reduce this?

OpenStudy (agent0smith):

I gave what you needed to reduce already...

OpenStudy (jkbo):

alright but the answer I got wasnt in my answer choices

OpenStudy (jkbo):

@agent0smith

OpenStudy (agent0smith):

The answer, 21/1150, is up there... it just doesn't look exactly like that.

OpenStudy (agent0smith):

For whatever reason, the answer is NOT in simplified form :/ but it's there.

OpenStudy (jkbo):

ok so it would be B? @agent0smith

OpenStudy (agent0smith):

If 21/1150 is the most simplified form (which it is), how can it then become 5/18 (which would be even more simplified than the most simplified form). The answer is not in the most simplified form. You could plug each answer into a calculator to see which matches 21/1150.

OpenStudy (jkbo):

o wait 1 goes int 21 so it would be D. @agent0smith

OpenStudy (agent0smith):

Again... the most simplified form, 21/1150, can NOT get any simpler. It can not become 1/48. Do exactly what i said in the last response (use a calculator) to find the answer.

OpenStudy (jkbo):

when I plugged in A. thats the only one that seemed to work

OpenStudy (agent0smith):

Check every answer on a calculator. Compare the results to 21/1150 on a calculator.

OpenStudy (jkbo):

can you give me an example I think Im doing it wrong then @agent0smith

OpenStudy (agent0smith):

you want me to give an example of how to use your calculator to divide? that might be difficult :P All you need to do is check each answer one at a time. One of them will look like this http://www.wolframalpha.com/input/?i=21%2F1150